The following table shows some data for three zero-coupon bonds. The face value of each bond is $1,000

Bond Price Maturity (Years) Yield to Maturity
A $ 380 20
B 380 10 %
C 18 9

a. What is the yield to maturity of bond A? (Do not round intermediate calculations. Enter your answer as a percent rounded to 3 decimal places. Assume annual compounding.)

b. What is the maturity of B? (Do not round intermediate calculations. Round your answer to 2 decimal places. Assume annual compounding.)

c. What is the price of C? (Do not round intermediate calculations. Round your answer to 2 decimal places. Assume annual compounding.)
Answer #1

a) Yield to maturity is calculated using the RATE function:-

=RATE(20,0,-380,1000)

=4.957%

b) Using NPER function:-

=NPER(10%,0,-380,1000)

=10.15

c) Using PV function:-

=PV(9%,18,0,1000)

=211.99